OpenStudy (anonymous):

write a polynomial function with rational coefficients in standard for whose zeros are x= -3,-4,2

OpenStudy (solomonzelman):

A polynomial function (with rational coefficients) with zeros: \(\large\color{black}{ \color{red}{a},~\color{green}{b},~\color{blue}{c} }\) is: \(\large\color{black}{ f(x)=(x- \color{red}{a})(x-\color{green}{b})(x-\color{blue}{c} }\))

OpenStudy (solomonzelman):

Do you want an example?

OpenStudy (anonymous):

so just plug in the numbers

OpenStudy (solomonzelman):

Yes, basically, your \(\large\color{black}{ \color{red}{a}}\) is \(\large\color{black}{ \color{red}{-3}, }\) your \(\large\color{black}{ \color{green}{b} }\) is \(\large\color{black}{ \color{green}{-4}, }\) and your \(\large\color{black}{ \color{blue}{c} }\) is \(\large\color{black}{ \color{blue}{2} . }\)

OpenStudy (solomonzelman):

After you plug in your numbers, you would then have to expand your function (I think).

OpenStudy (anonymous):

do u have the answer

OpenStudy (anonymous):

is it 24

OpenStudy (solomonzelman):

you are writing a function.... I don't see how would you get 24.

OpenStudy (anonymous):

Ok so u get f(x) = (x+3) (x+4) (x-1)

OpenStudy (solomonzelman):

you mean (x-2) at the end?

OpenStudy (anonymous):

well yes

OpenStudy (solomonzelman):

Good now you have to expand this, since you are asked to write a function with rational coefficients.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

oh ya

OpenStudy (solomonzelman):

I made a mistake.

OpenStudy (solomonzelman):

it is... \(\large\color{black}{ f(x)=(x+ 3)(x+4)(x-2 }\))

OpenStudy (solomonzelman):

So, you have to first start from expanding, \(\large\color{black}{ (x+ 3)(x+4) }\)

OpenStudy (anonymous):

ok i got tht

OpenStudy (solomonzelman):

sure, can you tell me what you got then, please?

OpenStudy (anonymous):

alright x squared + 7x +12

OpenStudy (solomonzelman):

yes, that is correct: So now instead of having, \(\large\color{black}{ f(x)=(x+ 3)(x+4)(x-2 ) }\) you have: \(\large\color{black}{ f(x)=(x^2+7x+12)(x-2 ) }\)

OpenStudy (solomonzelman):

you have to expand the last time and you are done.

OpenStudy (solomonzelman):

if you have a question regarding how to do it, please ask.

OpenStudy (anonymous):

ok i got x cubed + 5x squared -2x -24

OpenStudy (solomonzelman):

yes, that's right again: this is your function: \(\large\color{black}{f(x)=x^3+5x^2-2x-24 }\) (and just as asked, with rational coefficients....)

OpenStudy (solomonzelman):

Any more questions about the problem?

OpenStudy (solomonzelman):

btw, you can use my code: `\(\large\color{black}{ f(x)=x^3+5x^2-2x-24 }\)`

OpenStudy (anonymous):

no im good on tht one thank you but i have another

OpenStudy (solomonzelman):

yw;)

OpenStudy (anonymous):

\[^{3}\sqrt{a ^{16}}b ^{19}\] how to simplify

OpenStudy (anonymous):

anyone here

OpenStudy (solomonzelman):

\(\Large\color{black}{ \sqrt[3]{a^{16}b^{10}} }\) (made a size larger, `\(\Large\color{black}{ \sqrt[3]{a^{16}b^{10}} }\)` ) anyway, lets break it down like this: \(\Large\color{black}{ \sqrt[3]{a^{15}\times a\times b^{9}\times b} }\)

OpenStudy (anonymous):

alright then wut

OpenStudy (solomonzelman):

I lost my latex, got disconnected.

OpenStudy (anonymous):

ohh i thought you left me

OpenStudy (solomonzelman):

oh man, again~ I'll draw it, okay?

OpenStudy (anonymous):

alright

OpenStudy (solomonzelman):